OpenStudy (anonymous):

find and simplify the derivatives of f(t)=cosh^2(3t) - sinh^2(3t)

OpenStudy (anonymous):

trick question

OpenStudy (anonymous):

must be some more of that "trickonometry" that myininaya likes

OpenStudy (anonymous):

This should help: http://www.math.com/tables/derivatives/tableof.htm

OpenStudy (anonymous):

no it is a trick

myininaya (myininaya):

recall: \[\cosh(t)=\frac{e^t+e^{-t}}{2}\] and \[\sinh(t)=\frac{e^t-e^{-t}}{2}\]

OpenStudy (anonymous):

oh c'mon...

OpenStudy (anonymous):

...

myininaya (myininaya):

\[\cosh(3t)=\frac{e^{3t}+e^{-3t}}{2}\] and \[\sinh(3t)=\frac{e^{3t}-e^{-3t}}{2}\] -------------------------------------- \[([\cosh(3t)]^2)'=2\cosh(3t) \cdot \frac{3e^{3t}-3e^{-3t}}{2}=2\cdot \frac{e^{3t}+e^{-3t}}{2} \cdot \frac{3e^{3t}-3e^{-3t}}{2}\]

OpenStudy (anonymous):

what on earth. ok i guess i have to spill the beans

myininaya (myininaya):

i don't remember extra formulas

myininaya (myininaya):

theres no point

OpenStudy (anonymous):

supposed you were asked to find the derivative of \[\sin^2(3x)+\cos^2(3x)\] what would you get?

myininaya (myininaya):

yes i know the derivative of the thing is 0

OpenStudy (anonymous):

why you would get zero right? because the most basicest trig identity is \[\sin^2(x)+\cos^2(x)=1\]

myininaya (myininaya):

but wouldn't it be cooler to do all that algebra

OpenStudy (anonymous):

and the analogous one for hyperbolic functions is \[\cosh^2(x)-\sinh^2(x)=1\] so guess what the derivative is...

OpenStudy (anonymous):

so the answer is zero?

OpenStudy (anonymous):

cooler if you think it would be cooler to find the derivative of \[\sin^2(x)+\cos^2(x)=1\] by doing algebra